OpenStudy (anonymous):

Macy currently has a balance of $3,418.82 in an account earning simple interest. Eighteen years ago she opened the account with an initial deposit of $2,481. What is the interest rate on the account?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

So simple interest has the formula \[\large I = PRT\] Where I = interest earned P = principle balance R = rate (what we want) T = time in years

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

So we have T = 18 years P = the original 2481.00 R is what we want And to solve for I...since I = the interest earned...we need to do what she has in the account NOW - the original amount right? So first I = 3418.82 - 2481.00 After you find the I, plug it into the original formula and solve for R

OpenStudy (anonymous):

I= 937.82

OpenStudy (anonymous):

so then i plug that and i multiply?

OpenStudy (johnweldon1993):

Well yes now you plug it in \[\large 937.82 = 2481.00 \times R \times 18\] Which means \[\large R = \frac{937.82}{2481\times 18}\]

OpenStudy (anonymous):

thank you can you help me on another one @johnweldon1993